Find the volume of an object enclosed by cylinders z = y2 +2, z = 4−y2 and planes x = −1 and x = 2.



1
Expert's answer
2022-04-12T14:08:19-0400


Let's find the integration boundaries:

"-1\le x\le2" ;

to find the bounds for y equate "y^2 +2" and "4-y^2" :

"y^2+2=4-y^2"

"2y^2=2"

"y=\pm1"

"-1\le y\le1"

"y^2+2\le z\le 4-y^2"

"\displaystyle V=\int_{-1}^2dx\int_{-1}^1 dy\int_{y^2+2}^{4-y^2}dz=""\displaystyle \int_{-1}^2dx\int_{-1}^1 (4-y^2-(y^2+2))dy=""\displaystyle \int_{-1}^2dx\int_{-1}^1 (2-2y^2)dy=""\displaystyle \int_{-1}^2dx (2y-\frac23y^3)|_{-1}^1=\displaystyle \frac83\int_{-1}^2dx =""\displaystyle \frac83(2-(-1))=8"

Answer: "V=8".